OpenStudy (anonymous):

What are the solutions of x2 + 8x − 20 = 0?

OpenStudy (zzr0ck3r):

are you doing the quadratic formula, completing the square, or the AC method?

OpenStudy (zzr0ck3r):

Im guessing the AC method by your last question

OpenStudy (zzr0ck3r):

that was a question:)

OpenStudy (anonymous):

its the quadratic formula

OpenStudy (zzr0ck3r):

ok so (-b+-sqrt(b^2-4ac))/2a what are your a b and c?

OpenStudy (btaylor):

(x+10)(x-2) = 0 x+10 = 0 OR x-2=0 x = -10 or x = 2

OpenStudy (zzr0ck3r):

^ that is not the method they want you to do it with

OpenStudy (btaylor):

^ but it works. i don't care

OpenStudy (zzr0ck3r):

teachers dont accept that:) but good job

OpenStudy (anonymous):

that is exactly what i was looking for! Thank you @BTaylor

OpenStudy (zzr0ck3r):

a = 1 b = 8 and c = -20 so (-8+-sqrt(8^2-4(1)(-20))/2(1)

OpenStudy (btaylor):

yw

OpenStudy (zzr0ck3r):

ok so your not doing the quadratic formula?

OpenStudy (zzr0ck3r):

the point is to get to (x+10)(x-2) if you dont know how he got there i cant imagine your teacher will except that answer

OpenStudy (anonymous):

My teacher wanted the answer x=-10 x=2

OpenStudy (zzr0ck3r):

he does not want to see how you got there?

OpenStudy (zzr0ck3r):

he/she*

OpenStudy (anonymous):

its multiple choice.

OpenStudy (zzr0ck3r):

ahh, well you should understand how you got there:)

OpenStudy (anonymous):

i do understand

OpenStudy (zzr0ck3r):

great, seems strange to ask then:)